OpenStudy (sleepyjess):

Two triangles can be formed with the given information. Use the Law of Sines to solve the triangles. C = 67, a = 21, c = 20

OpenStudy (sleepyjess):

@satellite73

OpenStudy (anonymous):



OpenStudy (anonymous):

man that is one ugly picture, but i hope you get the idea

OpenStudy (sleepyjess):

Yep, I get it

OpenStudy (anonymous):

using the law of sines you get \[\frac{\sin(67)}{20}=\frac{\sin(A)}{21}\]

OpenStudy (anonymous):

when you take the arcsine you will get the acute A

OpenStudy (sleepyjess):

What would I take arcsin of? That is what is giving me trouble. :/

OpenStudy (anonymous):

you want angle A

OpenStudy (anonymous):

\[\sin(A)=\frac{21\sin(37)}{20}\] so that is what you need to find the arcsine of to find A

OpenStudy (anonymous):

but the arcsine obviously gives you only one answer the acute one